How can i solve this by the substitution method?
2M=N-3
3M=2N-9

Answer by checkley71(8403)

2M=N-3 OR M=(N-3)/2
SUBSTITUTING (N-3)/2 FOR M IN THE SECOND EQUATION WE GET 3(N-3)/2=2N-9 OR
(3N-9)/2=2N-9 CROSS MULTIPLYING WE GWT 2(2N-9)=3N-9 OR 4N-18=3N-9 OR
4N-3N=-9+18 OR N=9 AND M=(9-3)/2 OR M=6/2 OR M=3
PROOF 2(3)=9-3 OR 6=6 AND 3(3)=2(9)-9 OR 9=18-9 OR 9=9